OpenStudy (anonymous):

I WILL BECOME A FAN N GIVE U A MEDAL!!!!!<<<<< Jeff bought 10 cookies and ate 4 of them. He ate ____% of the cookies.

OpenStudy (igreen):

To find the percentage divide them and multiply by 100. 4 / 10 = 0.4 * 100 = 40%